For the function y = x2 - 6x + 8, perform the following tasks:
(a)Put the function in the form y = a(x-h)2 + k.
y = f(x)= ax2 + bx + c = 0
y = 1x2 + 6x + 8 = 0
a = 1 y = a(x-h)2 + k
b = 6 = a(1-6)2 + 8
c = 8 = a(5)2 + 8
= a(25)+ 8
= 33
(b)What is the equation for the line symmetry for the graph of this function?
(c)Graph the function using the equation in part a. Explain why it is not necessary to plot points to graph when using y=a(x-h)2+k.
(d)Describe how this graph compares to the graph of y=x2?

Answer by stanbon(75887)

For the function y = x^2 - 6x + 8, perform the following tasks:
(a)Put the function in the form y = a(x-h)^2 + k.
x^2-6x+?=y-8
x^2-6x+9=y-8+9
(x-3)^2=y+1
y=(x-3)^2-1
------------------
(b)What is the equation for the line symmetry for the graph of this function?
x=3
-----------------
(c)Graph the function using the equation in part a. Explain why it is not necessary to plot points to graph when using y=a(x-h)2+k.
Put the vertex at (3,-1) and draw the parabola up from there.
--------------------
(d)Describe how this graph compares to the graph of y=x^2?
The (x-3) moves the graph three to the right of the origin.
The -1 moves all the points down one.
Cheers,
Stan H.